OpenStudy (anonymous):

Meghan has 100 dollars more than Andrew has. After Meghan spends 20 dollars on groceries, she will have five times as much money as Andrew. How much money does Andrew have?

OpenStudy (anonymous):

let amount andrew has be x let amount meghan has be y therefore y = x + 100 ( first eqn ) now after meghan spends $20 , we have this eqn: (y-20) = 5x so y= x + 100 -- eqn 1 y= 5x + 20 -- eqn 2 equate then x + 100 = 5x +20 4x=80 x=20 but remember what we let x stand for. so andrew had $20